OpenStudy (anonymous):

Solve m plus one-fourth equals three-fourths. m equals negative one-half m equals one-fourth m equals one-half m equals one

OpenStudy (anonymous):

subtract 1/4 from 3/4 and what do you get?

OpenStudy (anonymous):

The denominator stays the same, so it's just 3 - 1

OpenStudy (anonymous):

So you get 2/4 which simplified is..?

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

no, divide 4 by 2 what is the answer?

OpenStudy (anonymous):

How many times does 2 go into 4?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

There you go, so then your answer is m = 1/2

OpenStudy (anonymous):

Choice C.

OpenStudy (anonymous):

Any other questions?

OpenStudy (anonymous):

The sum of a number and two-ninths is one and one-third. What is the number? one-ninth negative one-ninth one and one ninth negative one and one ninth

OpenStudy (anonymous):

Start by making 1 and 1/3 an improper fraction

OpenStudy (anonymous):

Actually, you don't need to do that

OpenStudy (anonymous):

we need to get a common denominator, so what do you need to do to 3 to get it to 9?

OpenStudy (anonymous):

3 times what is 9?

OpenStudy (anonymous):

27

OpenStudy (anonymous):

Not three times nine, we want our finishing number to be 9, so 3 x ? = 9?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

right, so since we multiplied the denominator of 1/3 by 3, we also need to multiply the numerator, so what is 1 x 3?

OpenStudy (anonymous):

So we then have 3/9. Here is our equation now: 2/9 + ? = 1 2/9

OpenStudy (anonymous):

whoops sorry typo: 2/9 + ? = 1 3/9

OpenStudy (anonymous):

NOW we want to make 1 3/9 an improper fraction so: 1 x 9 = 9 3 + 9 = 12 = 12/9

OpenStudy (anonymous):

Then we want to subtract: 12/9 - 2/9 = 10/9

OpenStudy (anonymous):

10/9 = 1 1/9

OpenStudy (anonymous):

So then we plug that into our equation and see if it works: 2/9 + 1 1/9 = 1 3/9 does this equation work?

OpenStudy (anonymous):

Hello? You there?

OpenStudy (anonymous):

The perimeter of a rectangular store receipt is ten and five-sixths inches. If three of the sides total eight and one-half inches, what is the length of the fourth side? two and one-sixth inches two and one-third inches 3 inches three and one-twelfth inches

OpenStudy (anonymous):

Did you get the last one?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Okay, first, what is the perimeter?

OpenStudy (anonymous):

Nvm, you just need to get a common denominator like I did in the last one

OpenStudy (anonymous):

find the common denominator of 10 5/6 and 8 1/2